OpenStudy (anonymous):

Hans bought a desktop computer and a laptop computer. Before finance charges, the laptop cost $450 more than the desktop. He paid for the computers using two different financing plans. For the desktop the interest rate was 9% per year, and for the laptop it was 6% per year. The total finance charges for one year were $300 . How much did each computer cost before finance charges?

OpenStudy (zpupster):

Let D = cost of the desktop Let L = cost of the laptop L = D + 450 Total interest rate paid for one year was 300.00 That interest rate was composed of D * .09 + L * .06 Formula is .09*D + .06*L = 300.00 Since L = D + 450, this formula becomes: .09*D + .06*(D + 450) = 300.00 Simplify this to get: .09*D + .06*D + .06*450 = 300.00 solve for D that will be the cost of desktop add 450 to get the laptop cost

OpenStudy (anonymous):

1820

OpenStudy (anonymous):

thanks

OpenStudy (zpupster):

yw